# Learning Conjugate Direction Fields for Planar Quadrilateral Mesh Generation

### v1

@@ 3 Preliminary
where $\mathbf{d}_{j,1},\mathbf{d}_{j,2}$ are a pair of unit principal direction vectors on the face $f_{j}$, and $\kappa_{j,1},\kappa_{j,2}$ are the corresponding principal curvatures.
@@
This flexibility, however, also presents a significant challenge: a CDF is ill-defined by geometry alone. To specify a single, desirable CDF from this high-dimensional space, additional conditions are required. Typically, these conditions are twofold:
@@
    Field Smoothness: To ensure a regular and coherent mesh layout while avoiding unnecessary singularities, the field must exhibit spatial smoothness across neighboring faces.
@@
    User-Defined Constraints: To capture design intent, the field must adhere to various user-specified conditions. A primary example is alignment with user-provided strokes on the surface.
@@
$\text{User-defined constraints}.$
@@
Figure 1: Visualization of direction fields and generated meshes on an ellipsoid.(a) Principal direction fields (PDF). (b) Conjugate direction fields (CDF). (c) PQ Mesh from PDF. (d) PQ Mesh from CDF.
@@ 4.4 Loss Functions
Note that our use of rotated ground truth vectors to measure alignment enables our formulation to handle the sign ambiguity for the ground-truth vectors, i.e., it allows the $\mathbf{u}_{j}$ to be close to either $\mathbf{u}^{*\perp}_{j}$ or $\mathbf{u}^{*\perp}_{j}$ (and the same for $\mathbf{v}_{j}$ and $\mathbf{v}^{*\perp}_{j}$). This is beneficial as both configurations are feasible for the subsequent parameterization and quad mesh generation.
However, for more effective learning, we shall not prescribe a specific correspondence between the predicted direction vectors $(\mathbf{u}_{j},\mathbf{v}_{j})$ and the ground truth directions $(\mathbf{u}_{j}^{*},\mathbf{v}_{j}^{*})$. Instead, what matters is that the predicted pair aligns well with the ground truth pair as a whole, regardless of which predicted vector corresponds to which ground truth vector. To handle this ambiguity, we also measure the alignment for the other correspondence with the following term:
@@
As discussed in Sec. 4.4, the specific correspondence between direction vectors $(\mathbf{u}_{j},\mathbf{v}_{j})$ and its neighboring direction vectors $(\mathbf{u}_{k},\mathbf{v}_{k})$ is ambiguous. Therefore, we also define the smoothness measure for the other correspondence as follows:
@@ 6.2 Computational Efficiency
Additionally, Fig. 5 illustrates the time complexity of our learning-based method over the traditional optimization-based method. For the two test shapes with increased face numbers, we can see a clear growth (approximately linear) of the optimization cost for CDF generation. In contrast, due to the highly parallelized feedforward process of our learning-based approach, the time cost increases only marginally. Fig. 6 compares the cost of CDF generation and quad mesh extraction for two test models shown in Fig. 3. We can see that for the optimization approach, CDF generation (even one round) takes the majority of the computational cost, and becomes a major bottleneck in the quad mesh design pipeline with iterative CDF attempts, especially for a bigger reference mesh such as Vase with $\sim$20k faces. On the contrary, our method can effectively address this issue and significantly reduce the time consumption, enabling a controllable and interactive CDF generation practice for user-friendly architectural design (see the supplementary video).
6.3 Result Quality

### v2

@@ 3 Preliminary
where $\mathbf{d}_{j,1},\mathbf{d}_{j,2}$ are unit principal direction vectors on $f_{j}$, and $\kappa_{j,1},\kappa_{j,2}$ are the corresponding principal curvatures.
@@
This flexibility, however, also presents a significant challenge: a CDF is ill-defined by geometry alone. To specify a single, desirable CDF from this high-dimensional space, additional conditions are required. Typical conditions include:
@@
    Field Smoothness: To ensure a regular and coherent mesh layout while avoiding unnecessary singularities, the field must be smooth across neighboring faces.
@@
    User-Defined Constraints: To capture design intent, the field should adhere to various user-specified conditions. One primary example is the alignment with user-provided strokes on the surface.
@@
$\text{User-specified constraints}.$
@@
Figure 1: Visualization of direction fields and generated meshes on an ellipsoid.
@@ 4.4 Loss Functions
Note that our use of rotated ground truth vectors to measure alignment enables us to handle the sign ambiguity for the ground-truth vectors, i.e., it allows the $\mathbf{u}_{j}$ to be close to either $\mathbf{u}^{*\perp}_{j}$ or $\mathbf{u}^{*\perp}_{j}$ (and the same for $\mathbf{v}_{j}$ and $\mathbf{v}^{*\perp}_{j}$). This is beneficial as both configurations are feasible for the subsequent parameterization and quad mesh generation.
However, for more effective learning, we shall not prescribe specific correspondence between the predicted direction vectors $(\mathbf{u}_{j},\mathbf{v}_{j})$ and the ground truth directions $(\mathbf{u}_{j}^{*},\mathbf{v}_{j}^{*})$. Instead, we only need the predicted pair to align well with the ground truth pair as a whole, regardless of which predicted vector corresponds to which ground truth vector. To handle this ambiguity, we also measure the alignment for the other correspondence with the following term:
@@
As discussed in Sec. 4.4, the correspondence between direction vectors $(\mathbf{u}_{j},\mathbf{v}_{j})$ and $(\mathbf{u}_{k},\mathbf{v}_{k})$ on neighboring faces is ambiguous. Therefore, we also define the smoothness measure for the other correspondence as follows:
@@ 6.2 Computational Efficiency
Additionally, Fig. 5 illustrates the time complexity of our learning-based method over the traditional optimization-based method. For two test shapes with increased face numbers, we can see a clear growth (approximately linear) of the optimization cost for CDF generation. In contrast, due to the highly parallelized feedforward process of our learning-based approach, the time cost increases only marginally. Fig. 6 compares the cost of CDF generation and quad mesh extraction for two test models shown in Fig. 3. We can see that for the optimization approach, CDF generation (even one round) takes the majority of the computational cost, and becomes a major bottleneck in the quad mesh design pipeline with iterative CDF attempts, especially for a bigger reference mesh such as Vase with $\sim$20k faces. On the contrary, our method can effectively address this issue and significantly reduce the time consumption, enabling a controllable and interactive CDF generation practice for user-friendly architectural design (see the supplementary video).
6.3 Result Quality